# Automated Label Unification for Multi-Dataset Semantic Segmentation with GNNs

### v1

@@ Abstract
Deep supervised models possess significant capability to assimilate extensive training data, thereby presenting an opportunity to enhance model performance through training on multiple datasets. However, conflicts arising from different label spaces among datasets may adversely affect model performance. In this paper, we propose a novel approach to automatically construct a unified label space across multiple datasets using graph neural networks. This enables semantic segmentation models to be trained simultaneously on multiple datasets, resulting in performance improvements. Unlike existing methods, our approach facilitates seamless training without the need for additional manual reannotation or taxonomy reconciliation. This significantly enhances the efficiency and effectiveness of multi-dataset segmentation model training. The results demonstrate that our method significantly outperforms other multi-dataset training methods when trained on seven datasets simultaneously, and achieves state-of-the-art performance on the WildDash 2 benchmark.
@@ 1 Introduction
Recent advances in computer vision [35, 21] have shown the benefits of large datasets for training robust visual models [2]. However, for deep supervised visual models that rely on annotated data, collecting large amounts of annotated data is costly [11]. To mitigate the high cost of annotation and utilize more data for training, several efforts [4, 23, 30] are underway to address the challenges of multi-dataset training, thereby utilizing multiple datasets to train more robust and generalized models.
Models trained on multiple datasets must confront the challenge of reconciling conflicting annotation standards and label spaces. For example, the class road in the BDD dataset [45] can be further divided into several classes in the Mapillary dataset [33]: road, lane marking and crosswalk. Similarly, the classes barrier and curb in the Mapillary dataset are labeled as curb in the IDD dataset [41]. These conflicts impact the supervised learning of models, as they may be incorrectly penalized for predicting finer-grained classes from other datasets.
Another challenge is the task of unifying diverse dataset labels to produce outputs in a standardized format [29, 24]. Several methods [11, 30] concatenate all datasets’ label spaces and then use language models to encode label names to construct text embedding space. There approaches introduce redundancy and fail to handle issues where labels with the same name have different annotation granularity. Other methods involve manually constructing universal taxonomies [4, 24] or relabeling [23], which can be time-consuming and labor-intensive. Recent approaches aim to automatically construct universal taxonomies [6, 40], but these methods typically identify inter-label relations between only two datasets, involving a time-consuming iterative training process.
@@
In this paper, we propose a novel approach leveraging Graph Neural Networks [20] to automatically construct unified label space, enabling segmentation model to be trained simultaneously on multiple datasets. In contrast to previous approaches [11, 4, 6], our approach eliminates the need for manual re-annotation or iterative training procedures to construct universal taxonomies, while also addressing the issues associated with language-based methods. As depicted in Figure 1, we utilize a language model to convert the dataset labels into text features. Then, we apply Graph Neural Networks (GNNs) to learn the relationships and associations among these labels. The process creates a unified label embedding space and dataset label mappings. The output head of the segmentation network incorporates the unified label embedding space to generate segmentation results in the unified label space. The dataset label mappings are subsequently utilized to align the unified segmentation results with the label spaces relevant to the dataset. This enables the training of segmentation models and graph neural networks with dataset annotations.
@@ 2 Related work
Figure 2: Illustration of our method that training with dataset-specific annotations through label mappings constructed by GNNs. We leverage a UniSegHead to enable simultaneous training on multiple datasets. In the UniSegHead, we compute the matrix product between pixel embedding and augmented unified node features output by the GNNs, resulting in predictions for the unified label space. We finally utilize the label mappings constructed by GNNs to map the unified predictions to dataset-specific prediction for training.
@@ 3.1 Unification of Multi-dataset Label Space
Unified Label Space. Given $K$ datasets with their respective label space $\{L_{1},L_{2},\ldots,L_{K}\}$, multi-dataset semantic segmentation requires a model to predict in a consistent label space that encompasses all dataset label spaces $L_{pred}=\bigcup_{i=1}^{K}L_{i}$. Each pixel needs to be assigned to a label in this unified label space. We define $N$ unified label nodes $\mathcal{A}=\{\alpha_{1},\alpha_{2},\ldots,\alpha_{N}\}$, serving as the nodes in the graph neural networks. Their corresponding $D$-dimensional learnable embedding $\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}\}$ represent the unified label embedding space. The number of unified label nodes $N$ is often smaller than the total number of dataset classes $|L|$, because we aim to merge multiple identical classes into one unified label. The image is first encoded into pixel embedding $\mathcal{P}$ by the segmentation network, which is then projected into the unified label embedding space to assign a unified label to each pixel.
Label Mappinngs. We define a mapping from the unified label space to the dataset-specific label space $\mathbf{M}_{i}:\mathcal{A}\rightarrow L_{i}$, which is used to train the model with dataset-specific annotations. Mathematically, $\mathbf{M}_{i}\in\{0,1\}^{N\times|L_{i}|}$ is a boolean linear transformation. Each unified class $\alpha$ is at most linked to a class $c$ within a specific dataset $i$ to prevent label conflicts: $\mathbf{M}_{\alpha}\mathbf{1}\leq\mathbf{1}$. To handle different annotation granularities, we use label mappings to merge multiple unified label nodes representing fine-grained classes into one super-class $c$: $\mathbf{M}_{c}^{\top}\mathbf{1}\geq\mathbf{1}$. For example, the curb from IDD can be simultaneously mapped by unified label nodes represented curb and barrier. We use unified label nodes as input nodes for the GNNs, which learn the label mappings and unified label embedding space, thus enabling the automated unification of multi-dataset label spaces.
@@ 3.2 Learning Unified Label Space with Graph Neural Networks
Input Nodes. To construct the input feature of dataset-specific label nodes, as illustrated in Figure 2, We first utilize ChatGPT [1] to generate richer textual descriptions for each label, providing the model with context to distinguish between different classes. Then, we employ llama-2 [39] to encode these label descriptions and generate text features. To further distinguish nodes from different datasets, we introduce learnable dataset embedding for each dataset $\{\mathbf{d}_{1},\mathbf{d}_{2},\ldots,\mathbf{d}_{K}\}$. The dataset embedding is combined with the text features to form the input features for dataset-specific label nodes:
@@
where $\mathbf{x}_{i,m}$ is the input feature of the $m$-th label from dataset $i$, and $f_{t}(l_{i,m})$ is the text feature of the label description $l_{i,m}$ encoded by language model. The input features of unified label nodes are randomly initialized to the same dimension as the dataset-specific label node. Together, dataset-specific label nodes and unified label nodes constitute the input nodes of GNNs.
Learnable Adjacency Matrix. To update the label mappings using gradient descent, we embed the label mappings as a continuous learnable graph adjacency matrix $\mathbf{M}_{a}$. Values in the adjacency matrix represent the weights of corresponding edges. Only the edges between unified label nodes and dataset-specific nodes are learnable, while others are fixed to zero. We apply softmax to the edges connecting unified label nodes with a specific dataset, ensuring that the sum of the edge weights $w$ between each unified label node and nodes of this dataset equals one. We formulate the element at the intersection of the $r$-th row and the $c$-th column in the lower triangular portion of $\mathbf{M}_{a}$ as:
@@
Unified Label Embedding. The forward propagation of our graph model follows the GraphSAGE framwork [17] as formulated in Equation 3. $\mathbf{W}^{k}$ and $\mathbf{X}^{k}$ are the weight and feature of the $k$-th GNNs layer. The $\sigma$ indicates nonlinear activation function such as tanh and ReLU. The last layer features of unified label nodes serve as the unified label embedding space $\mathbf{X}_{u}=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}]^{\top}$.
@@
To obtain the dataset label mappings, we partition the adjacency matrix into submatrices corresponding to each dataset. Each submatrix contains only the unified label nodes and the label nodes specific to that dataset. We compute the label mappings for each dataset based on the values in corresponding submatrix, divided into the following two cases. During training, we will alternately train the segmentation network and the GNNs . When training the GNNs, we directly use the values of the adjacency matrix as label mappings, facilitating updates using gradient descent algorithms. When training the segmentation network, we utilize the unbalanced optimal transport algorithm to solve for the boolean label mappings that satisfies the mapping constraints outlined in subsection 3.1. This algorithm detailed in Appendix C effectively handles the conversion of the continuous adjacency matrix into a discrete dataset label mappings required for training segmentation network.
@@ 3.3 Training a Universal Model with Dataset-specific Annotations
Training a universal model is divided into two steps. The first step involves training a robust encoder-decoder to provide the pixel embedding for each pixel position in the image, where similar objects should have similar features. The second step involves learning label mappings and unified label embedding space by GNNs. During the training phase, we alternate between these two steps, freezing one network while training another network. Both of these steps require supervised training using dataset-specific annotations. Here, we primarily focus on the training of GNNs, while further details of the training strategies are provided in the Appendix A.
Training with Dataset-specific Annotations. During training, an image is randomly sampled from dataset $i$ and fed to an segmentation network, which provides embedding for each pixel position $\mathcal{P}=\{\mathbf{p}_{1},\mathbf{p}_{2},\ldots,\mathbf{p}_{j}\}$, where $\mathbf{p}_{k}$ is $D$-dimensional vectors. We utilize a UniSegHead to assign a dataset label to each pixel. In the UniSegHead, We first project the pixel embedding into the unified label embedding space by multiplying the pixel embedding by the output feature of unified label node $\mathbf{X}_{u}$, as shown in Equation 4. Then, to train the universal model with dataset-specific annotations, we need to map the predictions in unified label space to dataset-specific label space to obtain the probabilities of dataset-specific classes. We could compute the per-pixel dataset-specific logits $\mathbf{s}$ by multiplying the dataset-specific label mappings $\mathbf{M}_{i}$ at each pixel:
@@
$\mathcal{L}_{ce}(\mathbf{y},\mathbf{s})=\sum^{L}_{i=1}y_{i}\log(\operatorname{softmax}(\mathbf{s})_{i}),$ (6)
@@
Orthogonality Loss. To achieve a conflict-free unified label space and avoid multiple unified label nodes representing the same class or excessively similar features, we impose soft constraints to promote orthogonality among the unified label node features, inspired by [38]. This orthogonality loss encourages unified label node embedding to be mutually orthogonal. It not only aligns the unified label nodes with annotation standards for practical use but also enhances the diversity of the model and helps in finding a better label mappings:
$\mathcal{L}_{orth}=\sum_{i=1}^{N}\operatorname{softmax}(\mathbf{X}_{u}\mathbf{x}_{i})_{i}\log(\operatorname{softmax}(\mathbf{X}_{u}\mathbf{x}_{i})_{i}).$ (7)
@@ 4 Experiments
Table 1: Training and test datasets in our experiments.
Driving scene CityScapes [13], Mapillary [33], BDD [45], IDD [41] WildDash [46], KITTI [16], CamVid [8],
Indoor scene SUN RGBD [36] ScanNet-20 [14]
@@
    Approach to construct unified label space. MC:Manually Construct, MR:Manually Relabel, DS:Dataset-specific, Auto:Automatically Construct.
@@ 4.1 Comparison on Multiple Datasets
In Table 2, we present the accuracy of our methods and compare them to other approaches on the seven training datasets. We use mean Intersection over Union (mIoU) to quantify the performance of models, a common metric used to evaluate the performance of segmentation models. Different methods adopt various approaches to construct their label spaces: Dataset-Specific represents a lack of a unified label space, where the model outputs a separate label space for each dataset. Manually Relabel means manually re-annotating each image. Manually Construct means the label space is constructed through human expertise. Automatically Construct includes methods where the unified label space is automatically constructed by the model. We also establish two baseline methods: Single dataset and Multi-SegHead. Single dataset demonstrates the results of training on individual dataset only, while Multi-SegHead trains on multiple datasets by using dataset-specific segmentation heads. The results demonstrate that our method achieves the best average performance in multi-dataset training, while also achieving significant performance improvements on datasets with a large number of classes such as the ADE and COCO datasets. We attribute this to the construction of a robust unified label space. Leveraging visual connections from the samples, our unified label space can discover label relationships beyond textual similarities. For instance, the visual appearance of the fireplace in ADE is similar to the tunnel in Mapillary. Despite their different semantic meanings, our model merges these labels for prediction, as detailed in subsection 4.5. This approach saves model capacity and facilitates knowledge transfer across datasets for improved prediction.
@@
In Table 3, as a supplement to Table 2, we compare our model with various models trained on single dataset, as well as each segmentation head output of the Multi-SegHead. Detailed data for each dataset can be found in the Appendix D. From the table, it can be observed that training with multiple datasets helps improve the model’s generalization performance. However, the performance of different segmentation head outputs in Multi-SegHead model shows significant differences due to the lack of a unified form of output that performs well across all datasets. In contrast, our approach provides a unified label space covering all datasets, resulting in a significant advantage in average performance. We also list the performance on the six unseen test datasets mentioned in Table 1. To evaluate the performance of our model on unseen datasets, we first evaluate the model results on its training dataset. We search for the optimal label mappings based on the accuracy of label predictions. There are no updates to any model parameters except the label mappings in this process. This process can actually be done manually without any annotation information. The results indicate that our model exhibits better generalization performance. It can handle various scenarios and consistently achieve excellent performance on the test set. Compared to Multi-SegHead model, our automatically constructed label space has advantages over dataset label spaces. The unified label space can integrate the semantic information from multiple dataset label spaces.
@@
Image Ground Truth BDD model MPL model ADE model Ours
@@
WildDash 2 provides a benchmark for semantic segmentation, designed to test the robustness of algorithms in real-world driving scenarios. Due to the insufficient number of training samples provided by this dataset, the official recommendation is to use multiple datasets for training. Therefore, this benchmark is well-suited to evaluate the effectiveness of multi-dataset training methods. The WildDash 2 dataset includes negative test cases to challenge the robustness of the model. These negative test cases mainly consist of unconventional driving scenarios, and even non-driving scenarios. Across all pixels within negative test images, a robust model is expected to predict the void label for open-set classes and anomalous objects. The WildDash 2 benchmark refers to the metric named Meta Avg mIoU Class, which calculates the mean Intersection over Union for each class by weighting negative and positive test cases according to their occurrence in the benchmark dataset.
Table 4 presents the current results on the WildDash 2 leaderboard. We provide results for zero-shot generalization and for our models trained on the CityScapes, IDD, BDD, Vistas and WildDash 2 datasets. The results indicate that our method achieves state-of-the-art performance in both zero-shot and trained settings. Our method exhibits significant performance improvements compared to other methods on negative test cases. We attribute this to the robustness of our model, which has been trained on diverse datasets, enabling it to perform well in unconventional scenarios. Our zero-shot model has been trained on a wider range of datasets compared to the trained model. Therefore, even without training on the WildDash 2 dataset, it achieves similar performance on negative test cases. Furthermore, our model predicts more classes compared to the 25 classes evaluated by WildDash 2. The zero-shot model predicts 231 classes, while the trained model predicts 70 classes. We map non-evaluated classes to a void label, enabling accurate predictions on negative test cases.
@@
Ours - 50.2 53.2 49.2 73.2 69.5 43.6
Table 5: Comparison of Different Variants of Construct Label Spaces.
Variants Label description Text encoder GNNs CS MPL SUN BDD IDD ADE COCO Mean
[…]
Ours 80.7 43.7 47.5 65.5 68.6 42.0 46.7 56.4

### v3

@@ Abstract
Deep supervised models possess significant capability to assimilate extensive training data, thereby presenting an opportunity to enhance model performance through training on multiple datasets. However, conflicts arising from different label spaces among datasets may adversely affect model performance. In this paper, we propose a novel approach to automatically construct a unified label space across multiple datasets using graph neural networks. This enables semantic segmentation models to be trained simultaneously on multiple datasets, resulting in performance improvements. Unlike existing methods, our approach facilitates seamless training without the need for additional manual reannotation or taxonomy reconciliation. This significantly enhances the efficiency and effectiveness of multi-dataset segmentation model training. The results demonstrate that our method significantly outperforms other multi-dataset training methods when trained on seven datasets simultaneously, and achieves state-of-the-art performance on the WildDash 2 benchmark. Our code can be found in https://github.com/Mrhonor/AutoUniSeg.
@@ 1 Introduction
Recent advances in computer vision [35, 21] have shown the advantages of large datasets in training robust visual models [2]. However, for deep supervised visual models that rely on annotated data, the collection of such extensive annotated datasets can be prohibitively costly [11]. To address this expense and expand the data available for training, several efforts [4, 23, 30] focus on the challenges of multi-dataset training, enabling the use of diverse datasets to train more robust and generalizable models.
Models trained on multiple datasets must confront the challenge of reconciling conflicting annotation standards and label spaces. For example, the class road in the BDD dataset [46] can be further divided into several classes in the Mapillary dataset [33]: road, lane marking and crosswalk. Similarly, the Mapillary dataset labels both barrier and curb as distinct classes, while in the IDD dataset [42], they are combined under the single label curb. These conflicts impact the supervised learning of models, as they may be incorrectly penalized for predicting finer-grained classes from other datasets.
Another challenge is the task of unifying diverse dataset labels to produce outputs in a standardized format [29, 24]. Several methods [11, 30] attempt to address this by concatenating the label spaces of all datasets and using language models to encode label names into a text embedding space. However, there approaches introduce redundancy and fail to handle issues where labels share names but differ in annotation granularity. Other methods involve manually constructing universal taxonomies [4, 24] or relabeling [23], both of which are time-consuming and labor-intensive. Recent approaches aim to automatically construct universal taxonomies [6, 41], but these methods typically identify inter-label relations between only two datasets, involving a time-consuming iterative training process.
@@
In this paper, we propose a novel approach leveraging Graph Neural Networks (GNNs) [20] to automatically construct unified label space, enabling segmentation model to be trained simultaneously on multiple datasets. In contrast to previous approaches [11, 4, 6], our approach eliminates the need for manual re-annotation or iterative training procedures to construct universal taxonomies, while also addressing the limitation of language-based methods in distinguishing categories with identical semantics. As depicted in Figure 1, we utilize a language model to convert the dataset labels into text features. Then, we apply GNNs to learn the relationships and associations among these labels. The process creates a unified label embedding space and dataset label mappings. The output head of the segmentation network incorporates the unified label embedding space to generate unified segmentation results within this unified space. The dataset label mappings are subsequently utilized to align the unified segmentation results with the label spaces relevant to each dataset. This enables the training of segmentation models and graph neural networks with dataset-specific annotations.
@@ 2 Related work
Figure 2: Illustration of our method that training with dataset-specific annotations through label mappings constructed by GNNs. We leverage a unified segmentation head (UniSegHead) to enable simultaneous training on multiple datasets. In the UniSegHead, we compute the matrix product between pixel embedding and augmented unified node features output by the GNNs, resulting in predictions for the unified label space. We finally utilize the label mappings constructed by GNNs to map the unified predictions to dataset-specific prediction for training.
@@ 3.1 Unification of Multi-dataset Label Space
Unified Label Space. Given $K$ datasets with their respective label space $\{L_{1},L_{2},\ldots,L_{K}\}$, multi-dataset semantic segmentation requires a model to predict within a consistent label space that encompasses all dataset label spaces $L_{pred}=\bigcup_{i=1}^{K}L_{i}$. Each pixel must be assigned to a label in this unified label space. We define $N$ unified label nodes $\mathcal{A}=\{\alpha_{1},\alpha_{2},\ldots,\alpha_{N}\}$, serving as the nodes in the graph neural networks. Their corresponding $D$-dimensional learnable embedding $\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}\}$ represent the unified label embedding space. The number of unified label nodes $N$ is often smaller than the total number of dataset classes $|L|$, because we aim to merge multiple identical classes into a single unified label. The image is first encoded into pixel embedding $\mathcal{P}$ by the segmentation network, which is then projected into the unified label embedding space to assign a unified label to each pixel.
Label Mappings. We define a mapping from the unified label space to the dataset-specific label space $\mathbf{M}_{i}:\mathcal{A}\rightarrow L_{i}$, which is used to train the model with dataset-specific annotations. Mathematically, $\mathbf{M}_{i}\in\{0,1\}^{N\times|L_{i}|}$ is a boolean linear transformation. Each unified class $\alpha$ is at most linked to a class $c$ within a specific dataset $i$ to prevent label conflicts: $\mathbf{M}_{\alpha}\mathbf{1}\leq\mathbf{1}$. To handle different annotation granularities, we use label mappings to merge multiple unified label nodes representing fine-grained classes into one super-class $c$: $\mathbf{M}_{c}^{\top}\mathbf{1}\geq\mathbf{1}$. For example, the curb from IDD can be simultaneously mapped by unified label nodes represented curb and barrier. We use unified label nodes as input nodes for the GNNs, which learn the label mappings and unified label embedding space, thus enabling the automated unification of multi-dataset label spaces.
@@ 3.2 Learning Unified Label Space with Graph Neural Networks
Input Nodes. To construct the input feature of dataset-specific label nodes, as illustrated in Figure 2, we used the dataset labels in the template "An image of <label> from the dataset <dataset>" as plain text input and employed ChatGPT to complete the detailed description of each label. Then, we employ llama-2 [40] to encode these label descriptions and generate text features. To further distinguish nodes from different datasets, we introduce learnable dataset embedding for each dataset $\{\mathbf{d}_{1},\mathbf{d}_{2},\ldots,\mathbf{d}_{K}\}$. The dataset embedding is combined with the text features to form the input features for dataset-specific label nodes:
@@
where $\mathbf{x}_{i,m}$ is the input feature of the $m$-th label from dataset $i$, and $f_{t}(l_{i,m})$ is the text feature of the label description $l_{i,m}$ encoded by language model. The input features of unified label nodes are randomly initialized to the same dimension as the dataset-specific label node. To determine the appropriate number of unified label nodes, inspired by the approach in [53], we used cross-validation results across different datasets to identify the number of mergeable categories, which served as the initial selection for the unified label nodes. The specific algorithm can be found in Appendix B. Together, dataset-specific label nodes and unified label nodes constitute the input nodes of GNNs. During the training process, we maintain a constant number of nodes. After the training is completed, we will remove inactive unified label nodes, meaning those that were not assigned to any dataset-specific label.
Learnable Adjacency Matrix. To enable label mappings to be updated via gradient descent, we embed the label mappings as a continuous, learnable graph adjacency matrix $\mathbf{M}_{a}$. Values in the adjacency matrix represent the weights of corresponding edges. Only the edges between unified label nodes and dataset-specific nodes are learnable, while others are fixed to zero. We apply a softmax operation to the edges connecting each unified label node with nodes of a particular dataset, ensuring that the sum of the edge weights $w$ between each unified label node and nodes of this dataset equals one. The element at the intersection of the $r$-th row and the $c$-th column in the lower triangular portion of $\mathbf{M}_{a}$ is formulated as:
@@
Unified Label Embedding. The forward propagation of our graph model follows the GraphSAGE framwork [17] as formulated in Equation 3. $\mathbf{W}^{k}$ and $\mathbf{X}^{k}$ are the weight and feature of the $k$-th GNNs layer. The $\sigma$ indicates nonlinear activation function, implemented as the tanh in this work. The output features of the unified label nodes from the final layer serve as the unified label embedding space, $\mathbf{X}_{u}=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}]^{\top}$.
@@
To obtain the dataset label mappings, we partition the adjacency matrix into submatrices, each corresponding to a specific dataset. Each submatrix contains only the unified label nodes and the label nodes specific to that dataset. We compute the label mappings for each dataset based on the values in its corresponding submatrix, divided into the following two cases. During training, we will alternately train the segmentation network and the GNNs. When training the GNNs, we directly use the value of the learnable adjacency matrix to establish label mappings, thereby facilitating weight updates through gradient descent. When training the segmentation network, we utilize the unbalanced optimal transport algorithm to solve for the boolean label mappings that satisfies the many-to-one mapping constraints. This algorithm detailed in Appendix C effectively handles the conversion of the continuous adjacency matrix into a discrete dataset label mappings required for training segmentation network.
@@ 3.3 Training a Universal Model with Dataset-specific Annotations
Training a universal model is divided into two steps. The first step involves training a robust encoder-decoder to provide the pixel embedding for each pixel position in the image, where similar objects should have similar features. The second step focuses on learning label mappings and unified label embedding space by GNNs. During the training phase, we alternate between these two steps, freezing one network while training another network. Both of these steps require supervised training using dataset-specific annotations. Here, we primarily focus on the training of GNNs, while further details of the training strategies are provided in the Appendix A.
Training with Dataset-specific Annotations. During training, an image is randomly sampled from dataset $i$ and fed into an segmentation network, which provides embedding for each pixel position $\mathcal{P}=\{\mathbf{p}_{1},\mathbf{p}_{2},\ldots,\mathbf{p}_{j}\}$, where $\mathbf{p}_{k}$ is $D$-dimensional vectors. We utilize a unified segmentation head (UniSegHead) to assign a dataset label to each pixel. In the UniSegHead, We first project the pixel embedding into the unified label embedding space by multiplying the pixel embedding by the output feature of unified label node $\mathbf{X}_{u}$, as shown in Equation 4. Then, to train the universal model with dataset-specific annotations, we need to map the predictions in unified label space to dataset-specific label space to obtain the probabilities of dataset-specific classes. This is achieved by computing per-pixel dataset-specific logits $\mathbf{s}$ by multiplying the dataset-specific label mappings $\mathbf{M}_{i}$ at each pixel:
@@
$\mathcal{L}_{ce}(\mathbf{y},\mathbf{s})=-\sum^{|L_{i}|}_{c=1}y_{c}\log(\operatorname{softmax}(\mathbf{s})_{c}),$ (6)
@@
Table 1: Training and test datasets in our experiments.
Driving scene CityScapes [13], Mapillary [33], BDD [46], IDD [42] WildDash 2 [47], KITTI [16], CamVid [8],
Indoor scene SUN RGBD [37] ScanNet-20 [14]
Orthogonality Loss. To achieve a conflict-free unified label space and avoid redundant unified label nodes that represent the same class or have overly similar features, we introduce soft constraints to promote orthogonality among the unified label node features, inspired by [39]. This orthogonality loss encourages unified label node embedding to be mutually orthogonal. It not only aligns the unified label nodes with annotation standards for practical use but also enhances the diversity of the model and helps in finding a better label mappings:
$\mathcal{L}_{orth}=-\sum_{i=1}^{N}\operatorname{softmax}(\mathbf{X}_{u}\mathbf{x}_{i})_{i}\log(\operatorname{softmax}(\mathbf{X}_{u}\mathbf{x}_{i})_{i}).$ (7)
@@ 4 Experiments
    Approach to construct label space. MC:Manually Construct, MR:Manually Relabel, DS:Dataset-specific, Auto:Automatically Construct.
@@ 4.1 Comparison on Multiple Datasets
In Table 2, we present the accuracy of our methods and compare them to other approaches on the seven training datasets. We use mean Intersection over Union (mIoU) to quantify the performance of models, a common metric used to evaluate the performance of segmentation models. Different methods adopt various approaches to construct their label spaces: Dataset-Specific represents a lack of a unified label space, where the model outputs a separate label space for each dataset. Manually Relabel means manually re-annotating each image. Manually Construct means the label space is constructed through human expertise. Automatically Construct includes methods where the unified label space is automatically constructed by the model. We also establish two baseline methods: Single dataset and Multi-SegHead. Single dataset demonstrates the results of training on individual dataset only, while Multi-SegHead trains on multiple datasets by using dataset-specific segmentation heads.
The results demonstrate that our method achieves the best average performance in multi-dataset training, while also achieving significant performance improvements on datasets with a large number of classes such as the ADE and COCO datasets. We attribute this to the construction of a robust unified label space. Leveraging visual connections from the samples, our unified label space can discover label relationships beyond textual similarities. For instance, the visual appearance of the fireplace in ADE is similar to the tunnel in Mapillary. Despite their different semantic meanings, our model merges these labels for prediction, as detailed in subsection 4.5. This approach saves model capacity and facilitates knowledge transfer across datasets for improved prediction.
@@
In Table 3, as a supplement to Table 2, we compare our model with various models trained on single dataset, as well as each segmentation head output of the Multi-SegHead. Detailed data for each dataset can be found in the Appendix D. From the table, it can be observed that training with multiple datasets helps improve the model’s generalization performance. However, the performance of different segmentation head outputs in Multi-SegHead model shows significant differences due to the lack of a unified form of output that performs well across all datasets. In contrast, our approach provides a unified label space covering all datasets, resulting in a significant advantage in average performance. We also list the performance on the five unseen test datasets mentioned in Table 1. To evaluate the performance of our model on unseen datasets, we first evaluate the model results on its training dataset. We search for the optimal label mappings based on the accuracy of label predictions. There are no updates to any model parameters except the label mappings in this process. This process can actually be done manually without any annotation information. The results indicate that our model exhibits better generalization performance. It can handle various scenarios and consistently achieve excellent performance on the test set. Compared to Multi-SegHead model, our automatically constructed label space has advantages over dataset label spaces. The unified label space can integrate the semantic information from multiple dataset label spaces.
@@
Image Ground Truth BDD Model MPL Model ADE Model Our 7ds Model
@@
WildDash 2 [47] provides a benchmark for semantic segmentation, designed to test the robustness of algorithms in real-world driving scenarios. Due to the insufficient number of training samples provided by this dataset, the official recommendation is to use multiple datasets for training. Therefore, this benchmark is well-suited to evaluate the effectiveness of multi-dataset training methods. The WildDash 2 dataset includes negative test cases to challenge the robustness of the model. These negative test cases mainly consist of unconventional driving scenarios, and even non-driving scenarios. Across all pixels within negative test images, a robust model is expected to predict the void label for open-set classes and anomalous objects. The WildDash 2 benchmark refers to the metric named Meta Avg mIoU Class, which calculates the mean Intersection over Union for each class by weighting negative and positive test cases according to their occurrence in the benchmark dataset.
Table 4 presents the current results on the WildDash 2 leaderboard. We present results for zero-shot generalization using our 7ds model. To evaluate in an unseen setting, we map the non-evaluated classes in the unified label space to a void label for both positive and negative test frames. To ensure a fair comparison with other works, we also provide evaluation results for models trained using the training datasets from the Robust Vision Challenge 2022[1], which include CityScapes, ADE20K, Vistas, VIPER [36], ScanNet, and WildDash 2. The results indicate that our method achieves state-of-the-art performance in both zero-shot and trained settings. Our method exhibits significant performance improvements compared to other methods on negative test cases. We attribute this to the robustness of our model, which has been trained on diverse datasets, enabling it to perform well in unconventional scenarios. Our zero-shot model has been trained on a wider range of datasets compared to the trained model. Therefore, even without training on the WildDash 2 dataset, it achieves similar performance on negative test cases.
@@
Ours - 50.0 52.2 47.5 72.4 68.6 44.6
Table 5: Comparison of Different Methods of Construct Label Spaces.
Methods $|L|$ CS MPL SUN BDD IDD ADE COCO Mean
Direct concatenation 448 79.2 38.8 46.9 64.3 66.7 33.3 38.2 52.5
Clustered by text features 329 79.9 40.8 47.3 65.6 68.6 33.8 38.0 53.4
Without GNN training 231 79.8 39.7 47.7 66.1 68.1 37.8 45.3 54.9
Without GPT’s label description 226 80.2 43.4 47.2 64.7 68.8 40.1 46.0 55.8
Our proposed method 217 80.7 43.7 47.5 65.5 68.6 42.0 46.7 56.4